# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ 2.1 Problem Formulation
Given an objective function $V$ and policy simplex $\Pi$, a policy $\pi\in\Pi$ is $\epsilon$-optimal if $V(\pi)\geq\sup_{\pi^{*}\in\Pi}V(\pi^{*})-\epsilon$.
@@
Heterogeneity among the local agents can be captured by modeling agent types as part of the agent state: assign a type $\varepsilon_{i}\in\mathcal{E}$ to each local agent $i\in[n]$ by letting $\mathcal{S}_{l}=\mathcal{E}\times\mathcal{S}^{\prime}_{l}$, where $\mathcal{E}$ is a set of possible types that are treated as a fixed part of the agent’s state. The transition and reward functions can vary depending on the agent’s type. The global agent can provide unique signals to local agents of each type by letting $s_{g}\in\mathcal{S}_{g}$ and $a_{g}\in\mathcal{A}_{g}$ denote a state/action vector where each element matches to a type $\varepsilon\in\mathcal{E}$.
2.2 Technical Background
Q-learning.
To provide background for the analysis in this paper, we review a few key technical concepts in RL. At the core of the standard Q-learning framework (Watkins & Dayan, 1992) for offline-RL is the $Q$-function $Q:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$. $Q$-learning seeks to produce a policy $\pi^{*}(\cdot|s)$ that maximizes the expected infinite horizon discounted reward. For any policy $\pi$, $Q^{\pi}(s,a)=\mathbb{E}^{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r(s(t),a(t))|s(0)=s,a(0)=a]$. One approach to learning the optimal policy $\pi^{*}(\cdot|s)$ is dynamic programming, where the $Q$-function is iteratively updated using value-iteration: $Q^{0}(s,a)=0$, for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. Then, for all $t\in[T]$, $Q^{t+1}(s,a)=\mathcal{T}Q^{t}(s,a)$, where $\mathcal{T}$ is the Bellman operator defined as
@@ 2.2 Technical Background
$\mathcal{T}Q^{t}(s,a)=r(s,a)+\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g}),\forall i\in[n]\end{subarray}}\max_{a^{\prime}\in\mathcal{A}_{g}\times\mathcal{A}_{l}^{n}}Q^{t}(s^{\prime},a^{\prime}).$ (4)
The Bellman operator $\mathcal{T}$ is $\gamma$-contractive, which ensures the existence of a unique fixed-point $Q^{*}$ such that $\mathcal{T}Q^{*}=Q^{*}$, by the Banach fixed-point theorem (Banach, 1922). Here, the optimal policy is the deterministic greedy policy $\pi^{*}:\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}\to\mathcal{A}_{g}\times\mathcal{A}_{l}^{n}$, where $\pi^{*}(s)=\arg\max_{a\in\mathcal{A}_{g}\times\mathcal{A}_{l}^{n}}Q^{*}(s,a)$. However, the complexity of a single update to the $Q$-function is $O(|\mathcal{S}_{g}||\mathcal{S}_{l}|^{n}|\mathcal{A}_{g}||\mathcal{A}_{l}|^{n})$, which grows exponentially with $n$. As the number of local agents increases $(n\boldsymbol{\mathit{g}}|\mathcal{S}_{l}|$), this exponential update complexity renders $Q$-learning impractical.
Mean-field Transformation.
To address this, mean-field MARL (under homogeneity assumptions) studies the distribution function $F_{z_{[n]}}:\mathcal{Z}_{l}\to\mathbb{R}$, where $\mathcal{Z}_{l}\coloneq\mathcal{S}_{l}\times\mathcal{A}_{l}$, given by
$F_{z_{[n]}}(z)\coloneq\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{s_{i}=z_{s},a_{i}=z_{a}\},\quad\forall z\coloneq(z_{s},z_{a})\in\mathcal{S}_{l}\times\mathcal{A}_{l}.$ (5)
Let $\mu_{n}(\mathcal{Z}_{l})=\{\frac{b}{n}|b\in\{0,\dots,n\}\}^{|\mathcal{S}_{l}|\times|\mathcal{A}_{l}|}$ be the space of $|\mathcal{S}_{l}|\times|\mathcal{A}_{l}|$-sized tables, where each entry is an element of $\{0,\frac{1}{n},\frac{2}{n},\dots,1\}$. In this space, $F_{z_{[n]}}\in\mu_{n}{(\mathcal{Z}_{l})}$ where $F_{z_{[n]}}$ represents the proportion of agents in each state/action pair. The $Q$-function is permutation-invariant in the local agents, since permuting the labels of homogeneous local agents with the same state will not change the action of the decision-making agent. Hence, $Q(s_{g},s_{[n]},a_{g},a_{[n]})=\widehat{Q}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})$. Here, $\widehat{Q}:\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mathcal{A}_{g}\times\mathcal{A}_{1}\times\mu_{n-1}(\mathcal{Z}_{l})\to\mathbb{R}$ is a reparameterized $Q$-function learned by mean-field value iteration: one initializes $\widehat{Q}^{0}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})=0$. At each time-step $t$, we update $\widehat{Q}$ as $\widehat{Q}^{t+1}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})=\widehat{\mathcal{T}}\widehat{Q}^{t}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})$, where $\widehat{\mathcal{T}}$ is the Bellman operator in distribution space:
$\widehat{\mathcal{T}}\widehat{Q}^{t}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})=r(s,a)+\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\forall i\in[n]\end{subarray}}\max_{\begin{subarray}{c}(a_{g}^{\prime},a_{1}^{\prime},a_{[n]\setminus 1}^{\prime})\\ \in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mathcal{A}_{l}^{n-1}\end{subarray}}\widehat{Q}^{t}(s^{\prime}_{g},s^{\prime}_{1},a^{\prime}_{g},a^{\prime}_{1},F_{z^{\prime}_{[n]\setminus 1}}).$ (6)
Since $\mathcal{T}$ is a $\gamma$-contraction, so is $\widehat{\mathcal{T}}$. Hence, $\widehat{T}$ has a unique fixed-point $\widehat{Q}^{*}$ such that $\widehat{Q}^{*}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})=Q^{*}(s_{g},s_{[n]},a_{g},a_{[n]})$ and the optimal policy is the deterministic greedy policy given by $\widehat{\pi}^{*}(s_{g},s_{1},F_{s_{[n]\setminus 1}})=\arg\max_{(a_{g},a_{1},a_{[n]\setminus 1})\in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mathcal{A}_{l}^{n-1}}\widehat{Q}^{*}(s_{g},s_{1},a_{g},a_{1},F_{s_{[n]\setminus 1},a_{[n]\setminus 1}},a_{g})$. The update complexity to the $\widehat{Q}$-function is $O(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{Z}_{l}|n^{|\mathcal{Z}_{l}|})$, which scales polynomially in $n$.

### v3

@@ 2.1 Problem formulation
Given a policy simplex $\Pi$, $\pi\in\Pi$ is an $\epsilon$-optimal policy if $V^{\pi}(s)\geq\sup_{\pi^{*}\in\Pi}V^{\pi^{*}}(s)-\epsilon$.
Motivating examples. Below we give examples of two cooperative MARL settings which are naturally modeled by our setting. Our experiments in Appendix B reveal a monotonic improvement in the learned policies as $k\to n$, while providing a substantial speedup over mean-field $Q$-learning.
Figure 1: Bounded exploration in warehouse accidents, and traffic congestions with Gaussian squeeze.
  • Gaussian squeeze: In this task, $n$ homogeneous agents determine individual actions $a_{i}$ to jointly maximize the objective $r(x)=xe^{-(x-\mu)^{2}/\sigma^{2}}$, where $x=\sum_{i=1}^{n}a_{i}$, and $\mu$ and $\sigma$ are the pre-defined mean and variance of the system. In scenarios of traffic congestion, each agent $i\in[n]$ is a controller trying to send $a_{i}$ vehicles into the main road, where controllers coordinate with each other to avoid congestion, hence avoiding either over-use or under-use, thereby contributing to the entire system. This GS problem is previously studied in Yang et al. (2018), and serves as an ablation study on the impact of subsampling for MARL.
  • Constrained exploration: Consider an $M\times M$ grid. Each agent’s state is a coordinate in $[M]\times[M]$. The state represents the center of a $d\times d$ box where the global agent constrains the local agents’ movements. Initially, all agents are in the same location. At each time-step, the local agents take actions $a_{i}(t)\in\mathbb{R}^{2}$ (e.g., up, down, left, right) to transition between states and collect rewards. The transition kernel ensures that local agents remain within the $d\times d$ box dictated by the global agent, using knowledge of $a_{i}(t),s_{g}(t)$, and $s_{i}(t)$. In warehouse settings where shelves have collapsed, creating hazardous or inaccessible areas, we want agents to clean these areas. However, exploration in these regions may be challenging due to physical constraints or safety concerns. Through an appropriate design of the reward and transition functions, the global agent could guide the local agents to focus on specific $d\times d$ grids, allowing efficient cleanup while minimizing risk.
Capturing heterogeneity.
Following Mondal et al. (2022); Xu and Klabjan (2023), our model can capture heterogeneity in the local agents by modeling agent types as part of the state: assign a type $\varepsilon_{i}\in\mathcal{E}$ to each local agent by letting $\mathcal{S}_{l}=\mathcal{E}\times\mathcal{S}^{\prime}_{l}$, where $\mathcal{E}$ is a set of possible types that are treated as a fixed part of the agent’s state. The transition and reward functions can vary depending on the agent’s type. The global agent can provide unique signals to local agents of each type by letting $s_{g}\in\mathcal{S}_{g}$ and $a_{g}\in\mathcal{A}_{g}$ denote a state/action vector where each element corresponds to a type.
3 Algorithmic Approach: Subsampled Value Iteration
Q-learning. Our starting point is the classic Q-learning framework (Watkins and Dayan, 1992) for offline-RL, which seeks to learn the $Q$-function $Q:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$. For any policy $\pi$, $Q^{\pi}(s,a)=\mathbb{E}^{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r(s(t),a(t))|s(0)\!=\!s,a(0)\!=\!a]$. Initially, $Q^{0}(s,a)\!=\!0$, for all $(s,a)\in\mathcal{S}\!\times\!\mathcal{A}$. Then, for all $t\!\in\![T]$, it is updated as $Q^{t+1}(s,a)\!\leftarrow\!\mathcal{T}Q^{t}(s,a)$, where the Bellman operator $\mathcal{T}$ is
$\mathcal{T}$ $Q^{t}(s,a)=r(s,a)+\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a),s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\forall i\in[n]\end{subarray}}\max_{a^{\prime}\in\mathcal{A}}Q^{t}(s^{\prime},a^{\prime}).$ (4)
It is well known that $\mathcal{T}$ is $\gamma$-contractive, ensuring that the above updates converge to a unique $Q^{*}$ such that $\mathcal{T}Q^{*}=Q^{*}$. The optimal policy $\pi^{*}\!:\!\mathcal{S}\!\to\!\mathcal{A}$ can then be computed greedily as $\pi^{*}(s)=\arg\max_{a\in\mathcal{A}}Q^{*}(s,a)$. However, the update complexity for the $Q$-function is $O(|\mathcal{S}||\mathcal{A}|)=O(|\mathcal{S}_{g}||\mathcal{S}_{l}|^{n}|\mathcal{A}_{g}||\mathcal{A}_{l}|^{n})$, which is exponential in the number of local agents increases.
Mean-field transformation. To address this, mean-field MARL (Yang et al., 2018) (under homogeneity assumptions) studies the empirical distribution function $F_{z_{[n]}}:\mathcal{Z}_{l}\to\mathbb{R}$, for $\mathcal{Z}_{l}\coloneq\mathcal{S}_{l}\times\mathcal{A}_{l}$:
$F_{z_{[n]}}(z)\coloneq\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{s_{i}=z_{s},a_{i}=z_{a}\},\quad\forall z\coloneqq(z_{s},z_{a})\in\mathcal{Z}_{l}\coloneqq\mathcal{S}_{l}\times\mathcal{A}_{l}.$ (5)
Let $\mu_{n}(\mathcal{Z}_{l})=\{\frac{b}{n}|b\in\{0,\dots,n\}\}^{|\mathcal{Z}_{l}|}$ be the space of $|\mathcal{Z}_{l}|$-sized vectors (or $|\mathcal{S}_{l}|\times|\mathcal{A}_{l}|$-sized tables), where each entry is in $\{0,{1}/{n},{2}/{n},\dots,1\}$. Intuitively, $\mu_{n}(\mathcal{Z}_{l})$ is a discrete distribution over $(\mathcal{S}_{l},\mathcal{A}_{l})$ where each probability assigned is a multiple of $1/n$. Then, $F_{z_{[n]}}\in\mu_{n}{(\mathcal{Z}_{l})}$ indicates the proportion of agents in each state-action pair.
As in Q-learning, in mean-field Q-learning initially $\hat{Q}^{0}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})=0$. At each time $t\in[T]$, we update $\hat{Q}$ as $\hat{Q}^{t+1}=\hat{\mathcal{T}}\hat{Q}^{t}$, where $\hat{\mathcal{T}}$ is the Bellman operator in distribution space:
$\hat{\mathcal{T}}\hat{Q}^{t}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})=r(s,a)+\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot\mid s_{g},a_{g})\\ s_{i}^{\prime}\sim P_{l}(\cdot\mid s_{i},s_{g},a_{i})\\ \forall i\in[n]\end{subarray}}\!\max_{\begin{subarray}{c}(a_{g}^{\prime},a_{1}^{\prime},a_{[n]\setminus 1}^{\prime})\\ \in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mathcal{A}_{l}^{n-1}\end{subarray}}\!\!\hat{Q}^{t}(s^{\prime}_{g},s^{\prime}_{1},a^{\prime}_{g},a^{\prime}_{1},F_{z^{\prime}_{[n]\setminus 1}})$
Since the $Q$-function is permutation-invariant in the homogeneous local agents, one sees that for each $t\geq 0$, $\smash{Q^{t}(s_{g},s_{[n]},a_{g},a_{[n]})=\hat{Q}^{t}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})}$. In other words, mean-field Q-learning implements the same updates as standard Q-learning. However, the update complexity of $\hat{Q}$ is only $\smash{O(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{Z}_{l}|n^{|\mathcal{Z}_{l}|})}$, which scales polynomially in $n$ but exponentially in $|\mathcal{Z}_{l}|$.
@@ 3 Algorithmic Approach: Subsampled Value Iteration
Existing methods use sample complexity $\tilde{O}(\min\{|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{Z}_{l}|^{n},|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{Z}_{l}|n^{|\mathcal{Z}_{l}|}\})$: one uses $Q$-learning if $|\mathcal{Z}_{l}|^{n-1}<n^{|\mathcal{Z}_{l}|}$, and mean-field value iteration otherwise. In each regime, as $n$ scales, the update complexity becomes computationally infeasible.
To further reduce the update complexity, we propose SUBSAMPLE-MFQ to overcome the polynomial (in $n$) sample complexity of mean-field $Q$-learning and the exponential (in $n$) sample complexity of traditional $Q$-learning. We begin by motivating the intuition behind our algorithms.
3.1 Overview of approach.
Offline Planning: Algorithm 1. First, the global agent randomly samples a subset of local agents $\Delta\subseteq[n]$ such that $|\Delta|=k$, for $k\leq n$. It then ignores all other local agents $[n]\setminus\Delta$, and performs value iteration (using $m$ samples to update the $Q$-function in each iterate) to approximately learn the $Q$-function $\hat{Q}_{k,m}^{\mathrm{est}}$ and policy $\hat{\pi}_{k,m}^{\mathrm{est}}$ for this surrogate subsystem of $k$ local agents. We denote the surrogate reward gained by this subsystem at each time step by $r_{\Delta}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, where
In Theorem E.3, we show that $\|\hat{Q}_{k,m}^{\mathrm{est}}-Q^{*}\|_{\infty}$ is Lipschitz continuous with respect to the TV-distance between the state/action pairs of the subsampled agents and the full set of $n$ agents. Equipped with this approximation guarantee, we show how to construct an approximately optimal policy on the full system on $n$ agents. In general, converting this policy on $k$ local agents to a policy on the full $n$-agent system without sacrificing error guarantees can be intractable, and there is a line on works on centralized-training decentralized-execution (CTDE) (Xu and Klabjan, 2023; Zhou et al., 2023) which shows that designing performant decentralized policies can be highly non-trivial.
Online Execution: Algorithm 2. In order to circumvent this obstacle and convert the optimality of each agent’s action in the $k$ local-agent subsystem to an approximate optimality guarantee on the full $n$-agent system, we propose a randomized policy ${\pi}^{\mathrm{est}}_{k,m}$, where the global agent samples $\smash{\Delta\in{[n]\choose k}}$ at each time-step to derive an action $\smash{a_{g}\leftarrow\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})}$, and where each local $i$ agent samples $\smash{k-1}$ other local agents $\Delta_{i}$ to derive an action $\smash{a_{i}\leftarrow\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta_{i}})}$. Finally, Theorem 4.4 shows that the policy ${\pi}_{k,m}^{\mathrm{est}}$ converges to the optimal policy $\pi^{*}$ as $\smash{k\to n}$ with rate $\tilde{O}(\frac{1}{\sqrt{k}})$.
3.2 Algorithm description.
We now formally describe the algorithms. Before describing Algorithm 1 (SUBSAMPLE-MFQ: Learning) and Algorithm 2 (SUBSAMPLE-MFQ: Execution) in detail, it will be helpful to first define the empirical distribution function:
Definition 3.2 (Empirical Distribution Function).
For any population $(z_{1},\dots,z_{n})\in\mathcal{Z}_{l}^{n}$, where $\mathcal{Z}_{l}\coloneq\mathcal{S}_{l}\times\mathcal{A}_{l}$, define the empirical distribution function $F_{z_{\Delta}}:\mathcal{Z}_{l}\to\mathbb{R}_{+}$ for all $z\coloneq(z_{s},z_{a})\in\mathcal{S}_{l}\times\mathcal{A}_{l}$ and for all $\smash{\Delta\in{[n]\choose k}}$ by $\smash{F_{z_{\Delta}}}(x)\coloneq\frac{1}{k}\sum_{i\in\Delta}\mathbf{1}\{s_{i}=z_{s},a_{i}=z_{a}\}$.
Let $\mu_{k}(\mathcal{Z}_{l})\coloneq\left\{\frac{b}{k}|b\in\{0,\dots,k\}\right\}^{|\mathcal{Z}_{l}|}$ be the space of $|\mathcal{Z}_{l}|$-sized vectors (or $|\mathcal{S}_{l}|\times|\mathcal{A}_{l}|$-sized tables) where each entry is in $\{0,1/k,2/k,\dots,1\}$. Intuitively, $\mu_{k}(\mathcal{Z}_{l})$ is a discrete distribution over $(\mathcal{S}_{l},\mathcal{A}_{l})$ where each probability assigned is a multiple of $1/k$. Here, $F_{z_{\Delta}}\in\mu_{k}(\mathcal{Z}_{l})$ indicates the proportion of agents (in the $k$-local-agent subsystem) at each state/action pair.
Algorithm 1 (Offline learning). Let $m\in\mathbb{N}$ denote the sample size for the learning algorithm with sampling parameter $k\leq n$. As in Remark 3.1, when $\smash{|\mathcal{Z}_{l}|^{k}\leq|\mathcal{Z}_{l}|k^{|\mathcal{Z}_{l}|}}$, the algorithm uses traditional value-iteration, and when $\smash{|\mathcal{Z}_{l}|^{k}>|\mathcal{Z}_{l}|k^{|\mathcal{Z}_{l}|}}$, it uses mean-field value iteration. We formally describe the procedure for each regime below:
Regime with Large State/Action Space: When $|\mathcal{Z}_{l}|^{k}\leq|\mathcal{Z}_{l}|k^{|\mathcal{Z}_{l}|}$, we iteratively learn the optimal $Q$-function for a subsystem with $k$-local agents denoted by $\smash{\hat{Q}_{k,m}^{t}:\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}\to\mathbb{R}}$, which is initialized to 0. At time $t$, we update $\hat{Q}_{k,m}^{t+1}(s_{g},{s_{\Delta}},a_{g},a_{\Delta})={\tilde{\mathcal{T}}}_{k,m}\hat{Q}_{k,m}^{t}(s_{g},s_{\Delta},a_{g},a_{\Delta})$, where ${\tilde{\mathcal{T}}}_{k,m}$ is the empirically adapted Bellman operator in Equation 7. Since the system is unknown, $\tilde{\mathcal{T}}_{k,m}$ cannot directly perform the Bellman update, so it instead uses $m$ random samples $s_{g}^{j}\sim P_{g}(\cdot|s_{g},a_{g})$ and $s_{i}^{j}\sim P_{l}(\cdot|s_{i},s_{g},a_{i})$ for each $j\in[m]$, $i\in\Delta$ to approximate the system:
@@
${\tilde{\mathcal{T}}}_{k,m}\hat{Q}_{k,m}^{t}(s_{g},{s_{\Delta}},a_{g},a_{\Delta})=r_{\Delta}(s,a)+\frac{\gamma}{m}\sum_{j\in[m]}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\hat{Q}_{k,m}^{t}(s_{g}^{j},{s_{\Delta}^{j}},a_{g}^{\prime},a_{\Delta}^{\prime}).$ (7)
Since $\tilde{\mathcal{T}}_{k,m}$ is $\gamma$-contractive, Algorithm 1 applies value iteration with $\tilde{\mathcal{T}}$ until $\hat{Q}_{k,m}$ converges to a fixed point satisfying $\smash{\tilde{\mathcal{T}}_{k,m}\hat{Q}_{k,m}^{\mathrm{est}}=\hat{Q}_{k,m}^{\mathrm{est}}}$, yielding a deterministic policy $\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}})$ where
$\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}})$ $\!=\!\!\mathop{\arg\max}_{a_{g}\in\mathcal{A}_{g},a_{\Delta}\in\mathcal{A}_{l}^{k}}\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}},a_{g},a_{\Delta}).$
Algorithm 1 SUB-SAMPLE-MFQ: Learning
0:  A multi-agent system as in Section 2, number of iterations $T$, sampling parameters $k\in[n]$, $m\in\mathbb{N}$, and discount factor $\gamma\in(0,1)$.
1:  Let $\Delta=\{1,\dots,k\}$, $\tilde{\Delta}=\{2,\dots,k\}$, and $\mu_{k-1}(\mathcal{Z}_{l})\!=\!\{\frac{b}{k-1}:b\in\{0,1,\dots,k-1\}\}^{|\mathcal{S}_{l}|\times|\mathcal{A}_{l}|}$.
2:  if $|\mathcal{Z}_{l}|^{k}\leq|\mathcal{Z}_{l}|k^{|\mathcal{Z}_{l}|}$ then
3:     // Sub-sampled version of standard $Q$-learning
4:     Initialize $\hat{Q}^{0}_{k,m}(s_{g},s_{\Delta},a_{g},a_{\Delta})=0,\forall(s_{g},s_{\Delta},a_{g},a_{\Delta})$.
5:     for $t=1$ to $T$ do
6:        for $(s_{g},s_{\Delta},a_{\Delta},a_{g})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$ do
7:           $\hat{Q}^{t+1}_{k,m}(s_{g},s_{\Delta},a_{g},a_{\Delta})\!=\!\tilde{\mathcal{T}}_{k,m}\hat{Q}^{t}_{k,m}(s_{g},s_{\Delta},a_{g},a_{\Delta})$
8:     Let the greedy policy be $\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})\coloneqq\arg\max_{\begin{subarray}{c}a_{g}\in\mathcal{A}_{g},a_{\Delta}\in\mathcal{A}_{l}^{k}\end{subarray}}\hat{Q}_{k,m}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})$.
9:  else
10:     // Sub-sampled version of mean-field $Q$-learning
11:     Initialize $\hat{Q}^{0}_{k,m}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g})=0$, $\forall(s_{g},s_{\Delta},a_{\Delta},a_{g})$.
12:     for $t=1$ to $T$ do
13:        for $(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{l}\times\mathcal{A}_{g}$ do
14:           $\hat{Q}^{t+1}_{k,m}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g})=\hat{\mathcal{T}}_{k,m}\hat{Q}^{t}_{k,m}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g})$
15:     Let the greedy policy be $\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\tilde{\Delta}}})\!:=\!\operatorname*{arg\,max}_{\begin{subarray}{c}a_{g}\in\mathcal{A}_{g},a_{i}\in\mathcal{A}_{l},\\ F_{a_{\tilde{\Delta}}}\in\mu_{k-1}{(\mathcal{A}_{l})}\end{subarray}}\!\!\hat{Q}_{k,m}^{T}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g})$.
Regime with Large Number of Agents: When $|\mathcal{Z}_{l}|^{k}>|\mathcal{Z}_{l}|k^{|\mathcal{Z}_{l}|}$, we learn the optimal mean-field $Q$-function for a $k$ local agent system, denoted by $\hat{Q}_{k,m}^{t}:\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{g}\times\mathcal{A}_{l}\to\mathbb{R}$, which is initialized to 0. At time $t$, we update $\hat{Q}_{k,m}^{t+1}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g})=\hat{\mathcal{T}}_{k,m}\hat{Q}_{k,m}^{t}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g})$, where $\hat{\mathcal{T}}_{k,m}$ is the empirically adapted mean-field Bellman operator in Equation 8. Similarly, as the system is unknown, $\hat{\mathcal{T}}_{k,m}$ cannot directly perform the Bellman update and instead uses $m$ random samples $s_{g}^{j}\sim P_{g}(\cdot|s_{g},a_{g})$ and $s_{i}^{j}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\forall j\in[m]$, $i\in\Delta$ to approximate the system:
$\!\!\!\hat{\mathcal{T}}_{k,m}\hat{Q}_{k,m}^{t}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g})\!=\!r_{\Delta}(s,a)\!+\!\frac{\gamma}{m}\!\sum_{j\in[m]}\!\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},a_{1}^{\prime}\in\mathcal{A}_{l},\\ F_{a_{\tilde{\Delta}}^{\prime}}\in\mu_{k-1}(\mathcal{A}_{l})\end{subarray}}\!\!\hat{Q}_{k,m}^{t}(s_{g}^{j},s_{1}^{j},F_{s_{\tilde{\Delta}}^{j},a_{\tilde{\Delta}^{\prime}}},a_{1}^{\prime},a_{g}^{\prime})$ (8)
Since $\hat{Q}_{k,m}^{t}$ depends on $s_{\Delta}$ and $a_{\Delta}$ through $F_{z_{\Delta}}$, $\hat{\mathcal{T}}_{k,m}$ is also $\gamma$-contractive. So, Algorithm 1 applies value iteration with $\hat{\mathcal{T}}$ until $\hat{Q}_{k,m}$ converges to a fixed point satisfying $\hat{\mathcal{T}}_{k,m}\hat{Q}_{k,m}^{\mathrm{est}}=\hat{Q}_{k,m}^{\mathrm{est}}$, yielding a deterministic policy $\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{1},F_{s_{\tilde{\Delta}}})$:
$\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{1},F_{s_{\tilde{\Delta}}})$ $\!=\!\!\mathop{\arg\max}_{\begin{subarray}{c}a_{g}\in\mathcal{A}_{g},a_{1}\in\mathcal{A}_{l},F_{a_{\tilde{\Delta}}}\in\mu_{k-1}(\mathcal{A}_{l})\end{subarray}}\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{1},F_{z_{\tilde{\Delta}}},a_{1},a_{g})$
For $\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta\setminus i})\!=\!a_{g}^{*},a_{i}^{*},a_{\Delta\setminus i}^{*}$, let $[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})]_{g}\!=\!a_{g}^{*}$ and $[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta\setminus i})]_{l}\!=\!a_{i}^{*}$.